Simplify
(-20^5b^13)/(5a^4b^14)

Answer by mrslong(1)

-20^5b^13 / 5a^4b^14 To simplify the following equation all you have to do is divide the whole numbers and subtract the superscripts from each other and which ever one has a value left over numerator or denominator you just place the value there.
-20^5 / 5 = -4^5 so the -4^5 is placed in the numerator
a^4 has nothing to subtract with or compare to so that value stays at the bottom.
b^13 /b^14 leaves you with -b at the bottom because 13 - 14 = (-1)
So your final answer would be -4^5/a^4-b